OpenStudy (kc_kennylau):

X~B(10,0.4) A pupil gets 2 correct, can the teacher conclude that the pupil has not studied well? (Using hypothesis and 5% confidence level)

OpenStudy (kc_kennylau):

Help me anyone? I suck at stats

OpenStudy (kc_kennylau):

@skullpatrol @Zarkon @ganeshie8 :/

OpenStudy (tkhunny):

This is a horrible question. If the student gets 10 correct, does that indicate good study habits? If the student gets zero correct, does that say ANYTHING about study habits? Maybe they were mugged on the way to the exam. Using the binomial distribution, what is the probability of exactly 2 correct out of 20, given B(10,0.40)?

OpenStudy (kc_kennylau):

out of 10

OpenStudy (tkhunny):

Right. I went back to fix that. What's the probability of making the same typo twice? \(p(0) = 0.6^{10}\) \(p(1) = 10*0.6^{9}*0.4\)

OpenStudy (kc_kennylau):

p(2)=45*0.6^8*0.4^2

OpenStudy (kc_kennylau):

Ah I have to go now :/

OpenStudy (tkhunny):

Now you have to decide if that is too small a probability to represent good study habits. You may wish to use the Normal Approximation to the Binomial.

OpenStudy (anonymous):

as u see, success rate is \[\left(\begin{matrix}10 \\ 2\end{matrix}\right) (0.4)^2(0.6)^8=\]?

OpenStudy (anonymous):

as then u will use it to predict the probability using basic stats and not taking in acct of other scenarios